OpenStudy (anonymous):

can anyone please help me or explain to me how to use the rational zero theorem???

OpenStudy (campbell_st):

ok... for rational zero theorem find the factors of the constant call them p find the factors of the coefficient of the leading term, call them q then a rational zero will be p/q as a simple example P(x) = 2x^2 +3x + 1 the factors of the constant are -1, 1 the factors of the coefficient of the leading term are -1, -2, 1, 2 so a rational zero exists at p/q so look at the possibles -1/1 and 1/-1 are the same so -1 1/-2, 1, 1/2 now test using P(-1), P(-1/2), P(1), P(1/2) and in doing so you see P(-1) = 0 P(-1/2) = 0 then rational zeros exist at x = -1 and x = -1/2 checking the polynomial can be factored. (2x + 1)(x + 1) = 0 x = -1/2, -1 thats a simple question but shows the process.

OpenStudy (anonymous):

how do you get :"the factors of the constant are -1, 1 the factors of the coefficient of the leading term are -1, -2, 1, 2" ?? @campbell_st

OpenStudy (anonymous):

this might help: http://www.sparknotes.com/math/algebra2/polynomials/section4.rhtml

OpenStudy (campbell_st):

what numbers multiply to 1 -1 x -1 = 1 and 1 x 1 = 1

OpenStudy (anonymous):

would you mind helping me out with this question @campbell_st ?? f(x)=2x^(4)+5x^(3)-8x^(2)-4x-7

OpenStudy (campbell_st):

ok... so the factors of 7 \[\pm 1, \pm 7\] factor of 2\[\pm 1, \pm 2\] so you need to look at P(1/2), P(7/2) P( -1/2) P(-7/2) P(7/1), P(-7/1), P(1/1) P(-1/1) I'd start with P(1) and P(-1) see if either are equal to zero.

OpenStudy (anonymous):

@jmprz_793 please check this out: http://openstudy.com/users/some_someone#/updates/5158c14ee4b07077e0c0a6aa it is a question previously asked by another user about the rational zero theorem. @campbell_st and I were explaining step by step.

OpenStudy (campbell_st):

I just pick up my calculator and solve it...

OpenStudy (campbell_st):

and you may want to check you have entered the equation correctly... as I only get 2 solutions. x = -3.54714 and x = 1.60699